The following table lists the components needed to assemble an end item, lead times (in weeks), and quantities on hand.

Item Lead Time Amount
on Hand
Direct Components
End 2 0 L(2), C(1), K(3)
L 2 8 B(2), J(3)
C 3 12 G(2), B(2)
K 3 17 H(4), B(2)
B 2 32
J 3 27
G 3 5
H 3 0


a. If 43 units of the end item are to be assembled, how many additional units of B are needed? (Hint: You don’t need to develop an MRP plan.)

Additional units            

b. An order for the end item is scheduled to be shipped at the start of week 13. What is the latest week that the order can be started and still be ready to ship on time? (Hint: You don’t need to develop an MRP plan.)

The latest week  
Answer #1

For producing 43 end units,

L units needed = 43*2 -8 = 78

C units needed = 43*1-12 =31

K units needed = 43*3-17 =112

additional Number of B units needed = 2*( 78+31+112)-32 = 410

2. For shipping order at start of week 13, it should start at the start if week 11.

For having L,C and K at start of week 11, they should start manufacturing L in week 9 and C,K in week 8.

B needs to start in week 6. J needs to start in week 6 while other components should start in week 5.

5 is the earliest week in which the work must start.